### v1

# How regular is the evolute of a plane curve?

Abstract.
We study the relationship between the smoothness of a plane curve and that of its evolute, especially in the cases where the parent curve is no more two or three times continuously differentiable, and exhibit the same kind of apparent improvement in regularity: in the generic local situation, the evolute has one order of regularity less than the parent curve.
Key words and phrases:
curvature, evolute, involute, regularity, cusps
2010 Mathematics Subject Classification:
53A04
The first named author was partially supported by the Bulgarian National Science Fund, Ministry of Education and Science of Bulgaria under contract KP-06-N82/6. The second named author wishes to thank the Institute of Mathematics and Informatics of the Bulgarian Academy of Sciences for its hospitality during the time when most of this work was carried out.
1. Introduction
1.1. Motivations and outline.
The notions of osculating circle and center of curvature of a smooth enough plane curve are classical, going back at least to Leibnitz. The locus of all the centers of curvature of a given curve is called its evolute and turns out to be the geometric envelope of the family of normal lines to the curve. The inverse operation–finding an involute, a curve which the given curve is the evolute of–has even found some applications to the design of some mechanisms, see for instance [11]. This has other beautiful geometrice properties, some less well-known, such as the Tait-Kneser Theorem [9]: if the radius of curvature varies in a monotone way, then the osculating circles are nested. Observe that the converse of this is immediate. Most of this can be found in elementary treaties about differential geometry or advanced calculus; a good succinct contemporary account can be found in [3], which singles out the Tait-Kneser Theorem as worthy of more attention.
However, the examples for which evolutes have been studied are mostly infinitely differentiable (cycloids), or even algebraic curves (conic sections), see the Wikipedia page [10] or e.g. [8]. When the curves are bounded, their curvature must admit at least four local extrema (Four Vertex Theorem [6]), which generate cusps in the evolute; the interesting generalization given in [2] (and subsequent papers) studies such singularities, but always in the context of the parent curve being infinitely differentiable.
In contrast to this, we are interested in the least possible amount of regularity to be required from the parent curve in order to get a certain regularity for the evolute, and conversely. At a first glance, it would seem that since the curvature depends on second derivatives and the evolute is expressed in terms of it, it should have two orders of regularity less than the parent curve. The result however is that the evolute has only one order of regularity less than the parent curve, provided that the curvature of that curve has non-vanishing derivative, see Theorem 1. This “bootstrap” phenomenon, an apparent regularity gain due to geometric constraints, reminds us of what happens when looking at distance functions and was pointed out in [5], who attribute the initial observation to the appendix of [4], see [7] for recent refinements.
Our results are local, since globally an evolute can have double points; we provide an example to that effect.
Much of the present paper may be known, but not always easy to find in current references, and we believe we have a relatively concise exposition of the results, and proofs which do not require many prerequisites.
1.2. Definitions and notations.
A plane curve is a continuous map $\gamma$ from an interval $(a;b)$ to the plane $\mathbb{R}^{2}$. For $k\in\mathbb{Z}_{+}$, $\alpha\in(0,1)$, we say the curve is of class $\mathcal{C}^{k}$ (resp. $\mathcal{C}^{k,\alpha}$) if its components are. In order to unify notation, we will write $\mathcal{C}^{k,0}$ for $\mathcal{C}^{k}$. We routinely use the fact that if $f,g\in\mathcal{C}^{k,\alpha}$, with $k\geq 1$, then $f\circ g\in\mathcal{C}^{k,\alpha}$ (this of course fails for $k=0,\alpha<1$).
For $k\geq 1$, we say that the curve is $\mathcal{C}^{k,\alpha}$-regular if in addition, for all $t\in(a;b)$, $\gamma^{\prime}(t)\neq(0,0)$. In this case, there exists an open interval $I$ containing $t$ such that $\gamma|_{I}$ embeds $I$ into $\mathbb{R}^{2}$ as a one-dimensional submanifold of the same regularity as $\gamma$. Note that this does not imply that $\gamma$ is globally one-to-one, only over a small enough interval.
For a curve of class $\mathcal{C}^{1}$ and $t_{0}\in(a;b)$, the arc length is $s_{\gamma}(t):=\int_{t_{0}}^{t}|\gamma^{\prime}(u)|du$. Changing the point $t_{0}$ changes $s_{\gamma}(t)$ by an additive constant, and we will have to do that on occasion. When $\gamma$ is, in addition, regular, $s_{\gamma}$ is an increasing diffeomorphism from $(a;b)$ to $(s_{\gamma}(a),s_{\gamma}(b))$, of the same smoothness class as $\gamma$. We will consider $\gamma\circ s_{\gamma}^{-1}$ as the “same curve parametrized by arc length”, but will need to distinguish it in notation.
When $\gamma$ is parametrized by arc length, $|\gamma^{\prime}(s)|=1$ for any $s\in(a;b)$. There is a lifting to $\mathbb{R}$ of the map $\gamma^{\prime}:(a;b)\longrightarrow S^{1}$, so we denote $\gamma^{\prime}(s)=(\cos\varphi(s),\sin\varphi(s))$ where $\varphi\in\mathcal{C}^{k-1,\alpha}(a;b)$ when $\gamma$ is of class $\mathcal{C}^{k,\alpha}$. Let $\nu(s):=(-\sin\varphi(s),\cos\varphi(s))$ (unit normal vector) be $\gamma^{\prime}(s)$ rotated by $\frac{\pi}{2}$.
When $k\geq 2$, $\kappa(s):=\varphi^{\prime}(s)$ is called the (signed) curvature of $\gamma$ and, if $\kappa(s)\neq 0$. Reducing the interval if needed, we make the standing assumption that for $t\in(a;b)$, $\kappa(t)\neq 0$. Then we can define $R(s):=1/\kappa(s)$ is the (signed) radius of curvature of $\gamma$. The well-known (plane) Frenet-Serret formulae immediately follow:
$\gamma^{\prime\prime}(s)=\frac{1}{R(s)}\nu(s),\qquad\nu^{\prime}(s)=-\frac{1}{R(s)}\gamma^{\prime}(s).$
The evolute of the curve $\gamma$ is given by the map
$\tilde{\gamma}(s):=\gamma(s)+R(s)\nu(s).$
When we parametrize the evolute by its own arc length $\tilde{s}$, we write $\tilde{s}\mapsto\tilde{\gamma}_{1}(\tilde{s})=\tilde{\gamma}(s)$.
1.3. Main results.
Theorem 1.
Let $\gamma:(a;b)\longrightarrow\mathbb{R}^{2}$ be a $\mathcal{C}^{2}$-regular curve such that $\kappa(t)\neq 0$ for all $t\in(a;b)$, and denote its evolute by $\tilde{\gamma}$. Then
  1. (i)
    $\tilde{\gamma}$ is $\mathcal{C}^{1}$-regular if and only if $R$ (or equivalently $\kappa$) is strictly monotone on $(a;b)$.
  2. (ii)
    For $k\geq 3$, $0\leq\alpha\leq 1$: $\gamma$ is $\mathcal{C}^{k,\alpha}$-regular and its radius of curvature has a nonzero derivative, if and only if $\tilde{\gamma}$ is $\mathcal{C}^{k-1,\alpha}$-regular.
Note that nonvanishing of the derivative of the radius of curvature implies its monotonicity, since we are working on an interval.
The fact that monotonicity of the radius of curvature is necessary and sufficient for smoothness of the evolute is well known, and mentioned for instance in [1, p. 308].
2. Proofs
Lemma 2.
Let $\gamma$ be a $\mathcal{C}^{2}$-regular plane curve parametrized over $(a,b)$ by arc length $s$, and $a<s_{1}<s_{2}<b$. Let $\tilde{\gamma}$ be its evolute, parametrized by the same $s\in(a,b)$. Then
(1) $\tilde{\gamma}(s_{2})-\tilde{\gamma}(s_{1})=\int_{s_{1}}^{s_{2}}\frac{R(s)-R(s_{2})}{R(s)}\gamma^{\prime}(s)ds+(R(s_{2})-R(s_{1}))\nu(s_{1}).$
Proof.
Observe first that the Frenet-Serret formulae (and the continuity of $R$) imply
$\nu(s_{2})-\nu(s_{1})=-\int_{s_{1}}^{s_{2}}\frac{1}{R(s)}\gamma^{\prime}(s)ds=-\frac{1}{R(s_{2})}(\gamma(s_{2})-\gamma(s_{1}))+\int_{s_{1}}^{s_{2}}\left(\frac{1}{R(s_{2})}-\frac{1}{R(s)}\right)\gamma^{\prime}(s)ds.$
Then notice that
$\tilde{\gamma}(s_{2})-\tilde{\gamma}(s_{1})=\gamma(s_{2})-\gamma(s_{1})+R(s_{2})(\nu(s_{2})-\nu(s_{1}))+(R(s_{2})-R(s_{1}))\nu(s_{1}).$
Proof of Theorem 1, (i).
Suppose that $R$ is strictly monotone on $(a;b)$, we may assume that it is increasing. Then, since $|\gamma^{\prime}(s)|=1$ and $0\leq R(s_{2})-R(s)\leq R(s_{2})-R(s_{1})$ for $s_{1}\leq s\leq s_{2}$,
$\left|\int_{s_{1}}^{s_{2}}\frac{R(s)-R(s_{2})}{R(s)}\gamma^{\prime}(s)ds\right|\leq|R(s_{2})-R(s_{1})|\int_{s_{1}}^{s_{2}}\frac{1}{|R(s)|}ds\leq|s_{2}-s_{1}|\frac{|R(s_{2})-R(s_{1})|}{\min_{[s_{1},s_{2}]}|R|}.$
So (1) implies
(2) $\frac{\tilde{\gamma}(s_{2})-\tilde{\gamma}(s_{1})}{R(s_{2})-R(s_{1})}=\nu(s_{1})+O(|s_{2}-s_{1}|).$
This implies that the curve admits a tangent vector, parallel to the unit vector $\nu(s)$, and that $s\mapsto R(s)$, which is a homeomorphism of intervals (by strict monotonicity) provides a (continuous) parametrization by arclength. Let $\tilde{s}:=R(s)$, $)\tilde{\gamma}_{1}(\tilde{s})=\tilde{\gamma}(R^{-1}(\tilde{s}$)) and $)\tilde{\gamma}_{1}^{\prime}(\tilde{s})=\nu(R^{-1}(\tilde{s}$)), which is a continuous non-vanishing vector function of $\tilde{s}$, so $\tilde{\gamma}_{1}$, and therefore $\tilde{\gamma}$, is $\mathcal{C}^{1}$-regular.
Conversely, assume that near a point $s_{1}$, $R$ is not strictly monotone.
Exchanging the roles of $s_{1}$ and $s_{2}$ in (1), we have
$\tilde{\gamma}(s_{2})-\tilde{\gamma}(s_{1})=\int_{s_{1}}^{s_{2}}\frac{R(s)-R(s_{1})}{R(s)}\gamma^{\prime}(s)ds+(R(s_{2})-R(s_{1}))\nu(s_{2}).$
First notice that we can find values of $s_{2}\geq s_{1}$ arbitrarily close to $s_{1}$ such that $|R(s_{2})-R(s_{1})|=\max_{s\in[s_{1};s_{2}]}|R(s)-R(s_{1})|$, and likewise for $s_{2}\leq s_{1}$. The argument above, applied to sequences of such values $s_{2,n}\to s_{1}$, leads to
$\frac{\tilde{\gamma}(s_{2,n})-\tilde{\gamma}(s_{1})}{R(s_{2,n})-R(s_{1})}=\nu(s_{2,n})+O(|s_{2,n}-s_{1}|),$
and since $\nu(s_{2,n})\to\nu(s_{1})$, in all cases $\nu(s_{1})$ belongs to the tangent cone of $\tilde{\gamma}$ at $\tilde{\gamma}(s_{1})$.
There are three cases.
Case 1.
If $R$ admits a strict local extremum at $s_{1}$, assume without loss of generality that it is a local maximum. With $\langle\cdot,\cdot\rangle$ standing for the Euclidean inner product in the plane,
(3) $\langle\tilde{\gamma}(s_{2})-\tilde{\gamma}(s_{1}),\nu(s_{1})\rangle=(R(s_{2})-R(s_{1}))+\int_{s_{1}}^{s_{2}}\frac{R(s)-R(s_{2})}{R(s)}\langle\gamma^{\prime}(s),\nu(s_{1})\rangle ds,$
so for any $\varepsilon>0$, if $s_{2}=s_{1}+h$ with $h$ chosen as above, $|R(s)-R(s_{2})|\leq|R(s_{1})-R(s_{2})|$ for $s\in[s_{1};s_{2}]$ and $\langle\gamma^{\prime}(s),\nu(s_{1})\rangle=\langle(\gamma^{\prime}(s)-\gamma^{\prime}(s_{1})),\nu(s_{1})\rangle=O(|s-s_{1}|)$, so the integral term is bounded in modulus by $C|s_{1}-s_{2}|^{2}|R(s_{1})-R(s_{2})|$ and $\langle\tilde{\gamma}(s_{2})-\tilde{\gamma}(s_{1}),\nu(s_{1})\rangle>0$. For values $h<0$ close enough to 0, the same argument again yields $\langle\tilde{\gamma}(s_{2})-\tilde{\gamma}(s_{1}),\nu(s_{1})\rangle>0$, so $\tilde{\gamma}$ cannot be regular at $s_{1}$.
Case 2.
If $R$ admits a non-strict local extremum at $s_{1}$, there must be values of $h$ arbitrarily close to 0 such that $R(s_{1}+h)=R(s_{1})$. So if $\tilde{\gamma}$ was differentiable at $s_{1}$, we would have $\tilde{\gamma}^{\prime}(s_{1})=0$, and this would be true for any (monotone) reparametrization of the curve, so it cannot be regular.
Case 3.
Finally if $R$ does not admits a strict local extremum at $s_{1}$, and is not strictly monotone in any neighborhood of $s_{1}$, there must be a sequence of values $h_{n}\to 0$ such that $R$ admits a local extremum at $s_{1}+h_{n}$, so $\tilde{\gamma}$ cannot be $\mathcal{C}^{1}$-regular at $s_{1}$.
Proof of Theorem 1, (ii).
Direct part:
Since $R(s)=\left(\langle\gamma^{\prime\prime}(s),\frac{\gamma^{\prime}(s)}{|\gamma^{\prime}(s)|}\rangle\right)^{-1}$, it is immediate that $R$ and $\tilde{\gamma}$ are of class $\mathcal{C}^{k-2,\alpha}$. The hypothesis about the derivative of $R$ means that $R$ is a $\mathcal{C}^{k-2,\alpha}$ diffeomorphism onto its image interval. When we reparametrize $\tilde{\gamma}$ by $R$, $)\tilde{\gamma}_{1}^{\prime}(\tilde{s})=\nu(R^{-1}(\tilde{s}$)). The vector $\nu$ is obtained by a rotation of $\pi/2$ from $\gamma^{\prime}(s)$ and so $s\mapsto\nu(s)$ is of class $\mathcal{C}^{k-1,\alpha}$. Since $R^{-1}$ is of class $\mathcal{C}^{k-2,\alpha}$ and $k-2\geq 1$, $\tilde{\gamma}_{1}^{\prime}$ is a $\mathcal{C}^{k-2,\alpha}$ non-vanishing vector function of $\tilde{s}$, so $\tilde{\gamma}_{1}$ is $\mathcal{C}^{k-1,\alpha}$-regular.
Converse:
Now assume that $\tilde{\gamma}$ is $\mathcal{C}^{k-1,\alpha}$-regular.
[…]
But Lemma 2 implies that $)\tilde{\gamma}_{1}^{\prime}(\tilde{s})=\nu(R^{-1}(\tilde{s}$)). The derivative $\frac{d\nu}{ds}$ never vanishes by the hypothesis that the curvature of the initial $\gamma$ is nonzero, and $R^{-1}$ is the inverse of a $\mathcal{C}^{1}$ function, so $|R^{-1}(\tilde{s}_{1})-R^{-1}(\tilde{s}_{2})|\geq C|s_{1}-s_{2}|$. If we had $(\nu\circ R^{-1})^{\prime}(\tilde{s}_{1})=0$, then we would have
$|\nu(R^{-1}(\tilde{s}_{2}))-\nu(R^{-1}(\tilde{s}_{1}))|=o(|s_{1}-s_{2}|)=o(|R^{-1}(\tilde{s}_{1})-R^{-1}(\tilde{s}_{2})|),$
so $\nu^{\prime}(R^{-1}(\tilde{s}_{1}))=0$, a contradiction.
Finally $\nu(s)=\tilde{\gamma}_{1}^{\prime}(R(s))$ must be of class $\mathcal{C}^{k-2,\alpha}$. Since $\gamma^{\prime}(s)$ is deduced from $\nu(s)$ by a rotation of angle $\frac{\pi}{2}$, $\gamma^{\prime}$ is of class $\mathcal{C}^{k-2,\alpha}$. Now the Frenet-Serret formulae give $\nu^{\prime}(s)=-\frac{1}{R(s)}\gamma^{\prime}(s)$, so $\nu^{\prime}$ is of class $\mathcal{C}^{k-2,\alpha}$ as a product of such. This implies that $\nu$ is of class $\mathcal{C}^{k-1,\alpha}$, so $\gamma^{\prime}$, deduced from it by a rotation of angle $\pi/2$, has the same smoothness, and $\gamma$ itself is of class $\mathcal{C}^{k,\alpha}$.
The fact that the derivative of $R$ does not vanish is equivalent to $\kappa^{\prime}\neq 0$, where $\kappa=1/R$ is the curvature of $\gamma$. The curvature of $\tilde{\gamma}$ is given by $\kappa^{3}/\kappa^{\prime}$; since we have assumed $\kappa(s)\neq 0$ for all $s$, and $\tilde{\gamma}$ is of class at least $\mathcal{C}^{2}$, the curvature of $\tilde{\gamma}$ is always bounded and $\kappa^{\prime}$, and thus $R^{\prime}$, do not vanish.
The fact that $R^{\prime}(s)\neq 0$ will also follow from the next section, which states more precise results about the behavior of the evolute in a neighborhood of a point $s_{1}$ where $\gamma$ is smooth and $R^{\prime}(s_{1})=0$. $\Box$
3. Local behavior
In this section, we will investigate the situations where $R^{\prime}(s)\neq 0$ no longer holds. We need to quantify the vanishing of the derivative of the (radius of) curvature.
Definition 3.
Let $I$ be an open interval, $f:I\longrightarrow R$ be a continuous function, and $s_{1}\in I$. We set
$m_{f}(s_{1})^{-}:=$ $\sup\{\mu\geq 0:\limsup_{s\to s_{1}}\frac{|f(s)-f(s_{1})|}{|s-s_{1}|^{\mu}}<\infty\}\in[0;\infty],$
$m_{f}(s_{1})^{+}:=$ $\inf\{\mu\geq 0:\liminf_{s\to s_{1}}\frac{|f(s)-f(s_{1})|}{|s-s_{1}|^{\mu}}>0\}\in[0;\infty].$
Note that $m_{f}(s_{1})^{+}\geq m_{f}(s_{1})^{-}$. When $f$ is of class $\mathcal{C}^{k}$, with $k\geq m_{f}(s_{1})$, and if there exists $l\in\{1,\dots,k\}$ such that $f^{(l)}(s_{1})\neq 0$,
$m_{f}(s_{1})=m_{f}(s_{1})^{+}=m_{f}(s_{1})^{-}\\ =\max\{m\geq 1:f^{(l)}(s_{1})=0,1\leq l\leq m\}+1=\min\{m\geq 1:f^{(m)}(s_{1})\neq 0\}.$
Notice that when this happens and when $f$ is strictly monotone on a neighborhood of $s_{1}$, $m_{f}(s_{1})$ must be an odd integer.
With the notations from the introduction, notably a curve $\gamma$ parametrized by arclength $s$, near a point $s_{1}$, the local frame for the evolute $\tilde{\gamma}$ will be given by the point $\tilde{\gamma}(s_{1})$ (which we will take to be the origin), and the direct orthonormal basis $(\nu(s_{1}),-\gamma^{\prime}(s_{1}))$. By a point $(x,y)$ we mean $\tilde{\gamma}(s_{1})+x\nu(s_{1})-y\gamma^{\prime}(s_{1})$.
Proposition 4.
Let $\gamma:(a;b)\longrightarrow\mathbb{R}^{2}$ be a $\mathcal{C}^{2}$-regular curve such that its signed radius of curvature $R$ is strictly monotone increasing on $(a;b)$, denote its evolute by $\tilde{\gamma}$. Let $s_{1}\in(a;b)$.
  1. (i)
    Suppose that there is a strictly increasing function $\omega:[0,\infty)\longrightarrow[0,\infty)$ with $\omega(0)=0$ such that for $s$ in a neighborhood of $s_{1}$, $|R(s)-R(s_{1})|\geq\omega(|s-s_{1}|)$. Then $\tilde{\gamma}$ can be represented in the local frame by $y=g(x)$, where $g(x)=O(x\omega^{-1}(x))$.
  2. (ii)
    If the evolute parametrized by its arclength, $\tilde{\gamma}_{1}$, is $\mathcal{C}^{1,\alpha}$-regular near $s_{1}$ for some $\alpha>0$, then $m_{R}(s_{1})^{-}\leq 1/\alpha$.
  3. (iii)
    In particular, if $\gamma$ is $\mathcal{C}^{k}$-regular and $1<m:=m_{R}(s_{1})\leq k$, $\tilde{\gamma}$ is $\mathcal{C}^{1,1/m}$-regular near $s_{1}$ with $m\geq 3$, and cannot be any smoother.
  4. (iv)
    If $\gamma$ is $\mathcal{C}^{k}$-regular and $m:=m_{R}(s_{1})$ is a nonzero even integer, then $R$ is not monotone in any neighborhood of $s_{1}$ and $\tilde{\gamma}$ admits a cusp: it is represented, changing $\nu$ into $\nu$ if needed, by the union of two graphs of the form $y=g_{1}(x)$ and $y=-g_{2}(x)$, $x\geq 0$, with $g_{1}(x),g_{2}(x)\geq 0$ and $g_{1}(x),g_{2}(x)\asymp x^{1+1/m}$.
Proof.
(i).
From (3) and the subsequent considerations we see that
(5) $\langle\tilde{\gamma}(s_{2})-\tilde{\gamma}(s_{1}),\nu(s_{1})\rangle=(R(s_{2})-R(s_{1}))+\int_{s_{1}}^{s_{2}}\frac{R(s)-R(s_{2})}{R(s)}\langle\gamma^{\prime}(s)-\gamma^{\prime}(s_{1}),\nu(s_{1})\rangle ds\\ =(R(s_{2})-R(s_{1}))(1+O(|s_{2}-s_{1}|)),\\ \langle\tilde{\gamma}(s_{2})-\tilde{\gamma}(s_{1}),\gamma^{\prime}(s_{1})\rangle=\int_{s_{1}}^{s_{2}}\frac{R(s)-R(s_{2})}{R(s)}\langle\gamma^{\prime}(s),\gamma^{\prime}(s_{1})\rangle ds\\ =O(|R(s_{2})-R(s_{1})||s_{2}-s_{1}|).$
This implies in particular that the projection from the image of $\tilde{\gamma}$ to the $x$-axis is one-to-one, therefore we can represent that image curve as a graph $y=g(x)$; and since $x\asymp R(s_{2})-R(s_{1})\geq\omega(s_{2}-s_{1})$ and $|y|\preceq|R(s_{2})-R(s_{1})||s_{2}-s_{1}|$, we have $|y|\preceq x\omega^{-1}(x)$.
(ii).
Since $\gamma$, parametrized by arclength, is (at least) $\mathcal{C}^{2}$-regular, we can write $s=\gamma^{-1}(\gamma(s))$, where $\gamma^{-1}$ is of class $\mathcal{C}^{2}$ and only defined locally on the embedded manifold $\gamma(I)$, $I$ a small enough interval. However $\gamma$ is also given as an involute of $\tilde{\gamma}$ by $\gamma_{1}(\tilde{s})=\tilde{\gamma}_{1}(\tilde{s})-\tilde{\gamma}_{1}^{\prime}(\tilde{s})(s-c)$, where $\tilde{s}$ is arclength on the evolute $\tilde{\gamma}$. By the regularity hypothesis, $\gamma_{1}$ will be of class $\mathcal{C}^{0,\alpha}$ and $)s=\gamma^{-1}(\gamma_{1}(\tilde{s}$)), so that there is a constant $C>0$ such that
$|s_{2}-s_{1}|\leq C|\tilde{s}_{2}-\tilde{s}_{1}|^{\alpha}=C|R(s_{2})-R(s_{1})|^{\alpha}\leq C_{\varepsilon}^{\prime}|s_{2}-s_{1}|^{\alpha(m_{R}(s_{1})^{-}-\varepsilon)}$
for any $\varepsilon>0$, so $\alpha m_{R}(s_{1})^{-}\leq 1$.
(iii).
The hypothesis implies that $R^{\prime}$ admits an isolated zero at $s_{1}$, so on a punctured neighborhood of $s_{1}$, $\tilde{\gamma}$ is of class at least $\mathcal{C}^{2}$ since $k\geq 3$. We can take $\omega(x)=Cx^{m}$ and deduce from (i) that $\tilde{\gamma}$ can be represented in the local frame by $y=g(x)$ et $g(x)=O(|x|^{1+1/m})$, in particular $\tilde{\gamma}$ is differentiable at $s_{1}$ with $\tilde{\gamma}_{1}^{\prime}(\tilde{s}_{1})=\nu(s_{1})$. More generally, $\tilde{\gamma}^{\prime}(R(s))=\nu(s)$ in a neighborhood of $s_{1}$ and
$|\nu(s_{2})-\nu(s_{1})|\preceq|s_{2}-s_{1}|\asymp|R(s_{2})-R(s_{1})|^{1/m},$
by the hypothesis on the order of $R$ at $s_{1}$.
The regularity cannot be improved as a consequence of (ii).
(iv).
Now $R$ has a strict local extremum. Suppose that it is a minimum, and that in a neighborhood of $s_{1}$ we have $R(s)-R(s_{1})=c(s-s_{1})^{m}+o(|s-s_{1}|^{m})$, with $c>0$. Then in the local frame, $0\leq x=R(s)-R(s_{1})\asymp(s-s_{1})^{m}$, and for $s\geq s_{1}$, $y\asymp(s-s_{1})^{m+1}$. For $s\leq s_{1}$, $y\asymp-(s-s_{1})^{m+1}$, and the claim easily follows. In the case of a local maximum, we change the sign of $\nu$ to obtain the same type of cusp in a modified local frame. ∎
End of Proof of Theorem 1 (ii), converse part.
We already know by the proof of Theorem 1 (i) that $R$ must be strictly monotone. If $R^{\prime}$ vanished at $s_{1}$, we would have to have $m_{R}(s_{1})\geq 3$, so $\tilde{\gamma}$ is at best of class $\mathcal{C}^{1,1/3}$ by Proposition 4(iii) and cannot be of class $\mathcal{C}^{2}$. $\Box$
4. Multiple points
Our definition of a regular curve restricts attention to a small enough interval in order to avoid problems with double points. Can those actually occur?
One can see that an evolute of a regular, one-to-one curve may exhibit double points.
Example 5.
There exists a $\mathcal{C}^{\infty}$ regular curve $\gamma$ such that its evolute is a limaçon $\tilde{\gamma}$ given, for instance, by the polar representation $r(\theta)=1+2\cos\theta$, $\frac{3\pi}{4}<\theta<\frac{3\pi}{4}$.
The curves $\gamma$ and $\tilde{\gamma}$ are represented in Figure 1.
Proof.
We start from a parametric representation for an arc of limaçon, which has a double point at $(0,0)$ corresponding to the values $\theta=-2\pi/3,2\pi/3$ of the parameter:
$\tilde{\gamma}(\theta)=\begin{pmatrix}{\cos\theta}\\ {\sin\theta}\end{pmatrix}(1+2\cos\theta).$
We deduce the arc length from $|\gamma^{\prime}(\theta)|=5+4\cos\theta$. An involute for $\tilde{\gamma}$ is given by
$\gamma(\theta)=\begin{pmatrix}{\cos\theta}\\ {\sin\theta}\end{pmatrix}\left(1+2\cos\theta+2(\sin\theta)\frac{\int_{-\pi}^{\theta}\sqrt{5+4\cos u}\,du}{\sqrt{5+4\cos\theta}}\right)\\ -\begin{pmatrix}{-\sin\theta}\\ {\cos\theta}\end{pmatrix}(1+2\cos\theta)\frac{\int_{-\pi}^{\theta}\sqrt{5+4\cos u}\,du}{\sqrt{5+4\cos\theta}}\\ =\begin{pmatrix}{\cos\theta}\\ {\sin\theta}\end{pmatrix}(1+2\cos\theta)+\begin{pmatrix}(\sin\theta)(1+4\cos\theta)\\ {-4\cos^{2}\theta-\cos\theta+2}\end{pmatrix}\frac{\int_{-\pi}^{\theta}\sqrt{5+4\cos u}\,du}{\sqrt{5+4\cos\theta}}.$
From the first formula, we easily see that $\gamma$ is regular. Notice that by what follows, $\gamma$ is also a simple curve. ∎
Figure 1.
Lemma 6.
The involute of a smooth curve, when it is a regular curve, has no double point.
[…]
On the other hand the same proof shows that if we start from a curve obtained as an evolute of a regular curve, the involute is regular outside of the possible points of contact with the evolute, which is not surprising since it will be a parallel to the parent curve.
Proof of Lemma 6.
By construction, the radius of curvature of an involute, being of the form $c-\tilde{s}$ must be strictly monotone. We conclude with the next result. $\Box$
Lemma 7.
A smooth regular curve with strictly monotone curvature cannot have a double point.
Proof.
Let us assume that with the parametrization we have chosen, the radius of curvature is strictly increasing. We claim that near each point of the curve, $\gamma(s)$ lies inside its osculating circle at $\gamma(s_{1})$ for $s<s_{1}$ and outside of it for $s>s_{1}$. Indeed the Tait-Kneser Theorem [9] tells us that osculating circles are nested within each other, so if $s_{1}<s_{2}$, the point $\gamma(s_{2})$ is on the osculating circle at $\gamma(s_{2})$, and $R(s_{2})>R(s_{1})$, so that circle is entirely outside the osculating circle at $\gamma(s_{1})$, on which lies $\gamma(s_{1})$. This way $\gamma(s_{2})$ can never become equal to a previously taken value. ∎

### v2

@@
 When the curves are bounded, their curvature must admit at least four local extrema (Four Vertex Theorem [6]), which generate cusps in the evolute; the interesting generalization given in [2] (and subsequent papers) studies such singularities, but always in the context of the parent curve being infinitely differentiable.
-In contrast to this, we are interested in the least possible amount of regularity to be required from the parent curve in order to get a certain regularity for the evolute, and conversely.
+In contrast to this, we are interested in the least possible amount of smoothness to be required from the parent curve in order to get a certain smoothness for the evolute, and conversely.
 At a first glance, it would seem that since the curvature depends on second derivatives and the evolute is expressed in terms of it, it should have two orders of regularity less than the parent curve.
@@
 In order to unify notation, we will write $\mathcal{C}^{k,0}$ for $\mathcal{C}^{k}$.
+Recall that a function is of class $\mathcal{C}^{\infty}$ if it is of class $\mathcal{C}^{k}$ for any $k$.
 We routinely use the fact that if $f,g\in\mathcal{C}^{k,\alpha}$, with $k\geq 1$, then $f\circ g\in\mathcal{C}^{k,\alpha}$ (this of course fails for $k=0,\alpha<1$).
@@
 Changing the point $t_{0}$ changes $s_{\gamma}(t)$ by an additive constant, and we will have to do that on occasion.
-When $\gamma$ is, in addition, regular, $s_{\gamma}$ is an increasing diffeomorphism from $(a;b)$ to $(s_{\gamma}(a),s_{\gamma}(b))$, of the same smoothness class as $\gamma$.
+When $\gamma$ is, in addition, regular, $s_{\gamma}$ is an increasing diffeomorphism from $(a;b)$ to $(s_{\gamma}(a),s_{\gamma}(b))$, of at least the same smoothness class as $\gamma$; the arclength parametrization therefore provides the best possible smoothness for the curve among the regular parametrizations.
 We will consider $\gamma\circ s_{\gamma}^{-1}$ as the “same curve parametrized by arc length”, but will need to distinguish it in notation.
@@
 Let $\nu(s):=(-\sin\varphi(s),\cos\varphi(s))$ (unit normal vector) be $\gamma^{\prime}(s)$ rotated by $\frac{\pi}{2}$.
-When $k\geq 2$, $\kappa(s):=\varphi^{\prime}(s)$ is called the (signed) curvature of $\gamma$ and, if $\kappa(s)\neq 0$.
-Reducing the interval if needed, we make the standing assumption that for $t\in(a;b)$, $\kappa(t)\neq 0$.
+When $k\geq 2$, $\kappa(s):=\varphi^{\prime}(s)$ is called the (signed) curvature of $\gamma$.
+Reducing the interval if needed, we make the standing assumption that for $s\in(a;b)$, $\kappa(s)\neq 0$.
 Then we can define $R(s):=1/\kappa(s)$ is the (signed) radius of curvature of $\gamma$.
@@
   2. (ii)
-    For $k\geq 3$, $0\leq\alpha\leq 1$: $\gamma$ is $\mathcal{C}^{k,\alpha}$-regular and its radius of curvature has a nonzero derivative, if and only if $\tilde{\gamma}$ is $\mathcal{C}^{k-1,\alpha}$-regular.
+    For $k\geq 3$, $0\leq\alpha\leq 1$: $\gamma$ is $\mathcal{C}^{k,\alpha}$-regular and $R^{\prime}(s)\neq 0$, if and only if $\tilde{\gamma}$ is $\mathcal{C}^{k-1,\alpha}$-regular.
 Note that nonvanishing of the derivative of the radius of curvature implies its monotonicity, since we are working on an interval.
@@
 Proposition 4.
-Let $\gamma:(a;b)\longrightarrow\mathbb{R}^{2}$ be a $\mathcal{C}^{2}$-regular curve such that its signed radius of curvature $R$ is strictly monotone increasing on $(a;b)$, denote its evolute by $\tilde{\gamma}$.
-Let $s_{1}\in(a;b)$.
+Let $\gamma:(a;b)\longrightarrow\mathbb{R}^{2}$ be a $\mathcal{C}^{2}$-regular curve, denote its evolute by $\tilde{\gamma}$. Let $s_{1}\in(a;b)$.
   1. (i)
-    Suppose that there is a strictly increasing function $\omega:[0,\infty)\longrightarrow[0,\infty)$ with $\omega(0)=0$ such that for $s$ in a neighborhood of $s_{1}$, $|R(s)-R(s_{1})|\geq\omega(|s-s_{1}|)$.
+    Suppose that the signed radius of curvature $R$ of $\gamma$ is strictly increasing on $(a;b)$ and let $\omega(x):=R(s_{1}+x)-R(s_{1})$.
 Then $\tilde{\gamma}$ can be represented in the local frame by $y=g(x)$, where $g(x)=O(x\omega^{-1}(x))$.
@@
     If the evolute parametrized by its arclength, $\tilde{\gamma}_{1}$, is $\mathcal{C}^{1,\alpha}$-regular near $s_{1}$ for some $\alpha>0$, then $m_{R}(s_{1})^{-}\leq 1/\alpha$.
+    If the derivative of $R$ vanishes to infinite order at $s_{1}$, the curve $\gamma$ is of class $\mathcal{C}^{1,0}$, and no better in the $\mathcal{C}^{1,\alpha}$ scale.
   3. (iii)
-    In particular, if $\gamma$ is $\mathcal{C}^{k}$-regular and $1<m:=m_{R}(s_{1})\leq k$, $\tilde{\gamma}$ is $\mathcal{C}^{1,1/m}$-regular near $s_{1}$ with $m\geq 3$, and cannot be any smoother.
+    In particular, if $\gamma$ is $\mathcal{C}^{k}$-regular, if $R$ is strictly increasing on $(a;b)$ and $1<m:=m_{R}(s_{1})\leq k$, $\tilde{\gamma}$ is $\mathcal{C}^{1,1/m}$-regular near $s_{1}$ with $m\geq 3$, and cannot be any smoother.
   4. (iv)
     If $\gamma$ is $\mathcal{C}^{k}$-regular and $m:=m_{R}(s_{1})$ is a nonzero even integer, then $R$ is not monotone in any neighborhood of $s_{1}$ and $\tilde{\gamma}$ admits a cusp: it is represented, changing $\nu$ into $\nu$ if needed, by the union of two graphs of the form $y=g_{1}(x)$ and $y=-g_{2}(x)$, $x\geq 0$, with $g_{1}(x),g_{2}(x)\geq 0$ and $g_{1}(x),g_{2}(x)\asymp x^{1+1/m}$.
+Notice that in item (i) we could have used any strictly increasing $\omega$ such that for $s$ in a neighborhood of $s_{1}$, $|R(s)-R(s_{1})|\geq\omega(|s-s_{1}|)$.
+This can apply to cases where $R$ is not monotone.
 Proof.
@@
 From (3) and the subsequent considerations we see that
-(5) $\langle\tilde{\gamma}(s_{2})-\tilde{\gamma}(s_{1}),\nu(s_{1})\rangle=(R(s_{2})-R(s_{1}))+\int_{s_{1}}^{s_{2}}\frac{R(s)-R(s_{2})}{R(s)}\langle\gamma^{\prime}(s)-\gamma^{\prime}(s_{1}),\nu(s_{1})\rangle ds\\ =(R(s_{2})-R(s_{1}))(1+O(|s_{2}-s_{1}|)),\\ \langle\tilde{\gamma}(s_{2})-\tilde{\gamma}(s_{1}),\gamma^{\prime}(s_{1})\rangle=\int_{s_{1}}^{s_{2}}\frac{R(s)-R(s_{2})}{R(s)}\langle\gamma^{\prime}(s),\gamma^{\prime}(s_{1})\rangle ds\\ =O(|R(s_{2})-R(s_{1})||s_{2}-s_{1}|).$
-This implies in particular that the projection from the image of $\tilde{\gamma}$ to the $x$-axis is one-to-one, therefore we can represent that image curve as a graph $y=g(x)$; and since $x\asymp R(s_{2})-R(s_{1})\geq\omega(s_{2}-s_{1})$ and $|y|\preceq|R(s_{2})-R(s_{1})||s_{2}-s_{1}|$, we have $|y|\preceq x\omega^{-1}(x)$.
+(5) $\langle\tilde{\gamma}(s_{2})-\tilde{\gamma}(s_{1}),\nu(s_{1})\rangle=(R(s_{2})-R(s_{1}))+\int_{s_{1}}^{s_{2}}\frac{R(s)-R(s_{2})}{R(s)}\langle\gamma^{\prime}(s)-\gamma^{\prime}(s_{1}),\nu(s_{1})\rangle ds\\ =(R(s_{2})-R(s_{1}))(1+O(|s_{2}-s_{1}|^{2})),\\ \langle\tilde{\gamma}(s_{2})-\tilde{\gamma}(s_{1}),\gamma^{\prime}(s_{1})\rangle=\int_{s_{1}}^{s_{2}}\frac{R(s)-R(s_{2})}{R(s)}\langle\gamma^{\prime}(s),\gamma^{\prime}(s_{1})\rangle ds\\ =O(|R(s_{2})-R(s_{1})||s_{2}-s_{1}|).$
+This implies in particular that the projection from the image of $\tilde{\gamma}$ to the $x$-axis is one-to-one, therefore we can represent that image curve as a graph $y=g(x)$; and since $x\asymp R(s_{2})-R(s_{1})$, $R(s_{2})-R(s_{1})\geq\omega(s_{2}-s_{1})$ for $s_{2}>s_{1}$, $R(s_{2})-R(s_{1})\leq-\omega(s_{2}-s_{1})$ for $s_{2}<s_{1}$, and $|y|\preceq|R(s_{2})-R(s_{1})||s_{2}-s_{1}|$, we have $|y|\preceq x\omega^{-1}(x)$.
 (ii).
@@
 However $\gamma$ is also given as an involute of $\tilde{\gamma}$ by $\gamma_{1}(\tilde{s})=\tilde{\gamma}_{1}(\tilde{s})-\tilde{\gamma}_{1}^{\prime}(\tilde{s})(s-c)$, where $\tilde{s}$ is arclength on the evolute $\tilde{\gamma}$.
-By the regularity hypothesis, $\gamma_{1}$ will be of class $\mathcal{C}^{0,\alpha}$ and $)s=\gamma^{-1}(\gamma_{1}(\tilde{s}$)), so that there is a constant $C>0$ such that
-$|s_{2}-s_{1}|\leq C|\tilde{s}_{2}-\tilde{s}_{1}|^{\alpha}=C|R(s_{2})-R(s_{1})|^{\alpha}\leq C_{\varepsilon}^{\prime}|s_{2}-s_{1}|^{\alpha(m_{R}(s_{1})^{-}-\varepsilon)}$
-for any $\varepsilon>0$, so $\alpha m_{R}(s_{1})^{-}\leq 1$.
+By the regularity hypothesis, $\gamma_{1}$ will be of class $\mathcal{C}^{0,\alpha}$ and $)s=\gamma^{-1}(\gamma_{1}(\tilde{s}$)), so there is a constant $C>0$ such that
+$|s_{2}-s_{1}|\leq C|\tilde{s}_{2}-\tilde{s}_{1}|^{\alpha}=C|R(s_{2})-R(s_{1})|^{\alpha};$
+if $m_{R}(s_{1})^{-}<\infty$, for any $\varepsilon>0$ we have $C_{\varepsilon}$ such that $|R(s_{2})-R(s_{1})|\leq C_{\varepsilon}|s_{2}-s_{1}|^{m_{R}(s_{1})^{-}-\varepsilon}$, so that $|s_{2}-s_{1}|\leq C_{\varepsilon}^{\prime}|s_{2}-s_{1}|^{\alpha(m_{R}(s_{1})^{-}-\varepsilon)}$, so $\alpha m_{R}(s_{1})^{-}\leq 1$.
+If $m_{R}(s_{1})^{-}=\infty$, $|R(s_{2})-R(s_{1})|\leq C|s_{2}-s_{1}|^{1+1/\alpha}$, which leads to a contradiction.
+For the second statement, Theorem 1 implies that $R$ is of class $\mathcal{C}^{1}$.
+If $R$ was of class $\mathcal{C}^{1,\alpha}$, $\alpha>0$, we would have $m_{R}(s_{1})^{-}\leq 1/\alpha$ which contradicts the hypothesis of vanishing of the derivative to infinite order.
 (iii).
 The hypothesis implies that $R^{\prime}$ admits an isolated zero at $s_{1}$, so on a punctured neighborhood of $s_{1}$, $\tilde{\gamma}$ is of class at least $\mathcal{C}^{2}$ since $k\geq 3$.
-We can take $\omega(x)=Cx^{m}$ and deduce from (i) that $\tilde{\gamma}$ can be represented in the local frame by $y=g(x)$ et $g(x)=O(|x|^{1+1/m})$, in particular $\tilde{\gamma}$ is differentiable at $s_{1}$ with $\tilde{\gamma}_{1}^{\prime}(\tilde{s}_{1})=\nu(s_{1})$.
-More generally, $\tilde{\gamma}^{\prime}(R(s))=\nu(s)$ in a neighborhood of $s_{1}$ and
-$|\nu(s_{2})-\nu(s_{1})|\preceq|s_{2}-s_{1}|\asymp|R(s_{2})-R(s_{1})|^{1/m},$
-by the hypothesis on the order of $R$ at $s_{1}$.
+We have $\omega(x)\asymp x^{m}$ and deduce from (i) that $\tilde{\gamma}$ can be represented in the local frame by $y=g(x)$ et $g(x)=O(|x|^{1+1/m})$, in particular $\tilde{\gamma}$ is differentiable at $s_{1}$ with $\tilde{\gamma}_{1}^{\prime}(\tilde{s}_{1})=\nu(s_{1})$.
+More generally, for $s,s^{\prime}$ in a neighborhood of $s_{1}$, $\tilde{\gamma}^{\prime}(R(s))=\nu(s)$ and $|\nu(s^{\prime})-\nu(s_{1})|\preceq|s^{\prime}-s|$; we claim that
+$|s^{\prime}-s|\preceq|R(s^{\prime})-R(s)|^{1/m}.$
+To see this, assume without loss of generality that $R^{(m)}(t)\geq c>0$ for $t$ in an open interval $I$ around $s_{1}$, and $s,s^{\prime}\in I$, $s_{1}\leq s<s^{\prime}$.
+An easy induction argument then shows that for $0\leq j\leq m-1$, $R^{(m-j)}(s)\geq\frac{c}{j!}(s-s_{1})^{j}\geq 0$.
+By the Taylor-Lagrange formula with integral remainder,
+$R(s^{\prime})-R(s)=\sum_{j=1}^{m-1}\frac{(s^{\prime}-s)^{j}}{j!}R^{(j)}(s)+\int_{s}^{s^{\prime}}\frac{(s^{\prime}-t)^{m-1}}{(m-1)!}R^{(m)}(t)dt\\ \geq c\int_{s}^{s^{\prime}}\frac{(s^{\prime}-t)^{m-1}}{(m-1)!}dt=c\frac{(s^{\prime}-s)^{m}}{m!}.$
+This proves the claim.
 The regularity cannot be improved as a consequence of (ii).
@@
 We already know by the proof of Theorem 1 (i) that $R$ must be strictly monotone.
-If $R^{\prime}$ vanished at $s_{1}$, we would have to have $m_{R}(s_{1})\geq 3$, so $\tilde{\gamma}$ is at best of class $\mathcal{C}^{1,1/3}$ by Proposition 4(iii) and cannot be of class $\mathcal{C}^{2}$.
+If $R^{\prime}$ vanished at $s_{1}$, we would have to have $m_{R}(s_{1})\geq 3$, so $\tilde{\gamma}$ is at best of class $\mathcal{C}^{1,1/3}$ by Proposition 4(ii) and (iii), and cannot be of class $\mathcal{C}^{2}$.
 $\Box$